The coordinates of the vertices of a triangle are A( x1, y1), B( x2, y2), and C(x3,y3) .Then what's the co ordinates of the circumcentre.
1
Expert's answer
2020-10-26T20:37:00-0400

"A(x_1,y_1), \ B(x_2,y_2),\ C(x_3,y_3)"

To calculate the coordinate of the circumcenter, we have to solve any 2 bisector equations of any of the three lines AB, BC, or AC, then find out the intersection points. i.e.



The midpoint of side AB = "\lgroup \dfrac{x_1 + x_2}{2}, \dfrac{y_1 + y_2}{2} \rgroup= (x_{12}, y_{12})"


And the slope of AB = "\dfrac{y_2-y_1}{x_2-x_1} = m_{12}"

"\therefore" The slope of the perpendicular bisector of side AB = "\frac{-1}{m_{12}} = m_{21}"


The equation of the perpendicular bisector of AB is therefore,

"y-y_{12} = m_{21}(x-x_{12}) \ ---(i)"


Similarly,

For side AC

The midpoint of side AC =

"\lgroup \dfrac{x_1 + x_3}{2}, \dfrac{y_1 + y_3}{2} \rgroup= (x_{13}, y_{13})"


And its slope = "\dfrac{y_3-y_1}{x_3-x_1} = m_{13}"

Making the slope of its perpendicular bisector = "\frac{-1}{m_{13}} = m_{31}"


The equation of the perpendicular bisector of length AC is therefore,

"y-y_{13} = m_{31}(x-x_{13}) \ ---(ii)"


On solving equations (i) and (ii), the values of x and y is the coordinate of the circumcenter


OR, using formula method;


"Circumcenter(X, Y) = \lgroup \dfrac{x_1 sin 2A + x_2 sin 2B + x_3 sin 2C}{sin 2A + sin 2B + sin 2C},\dfrac{y_1 sin 2A + y_2 sin 2B + y_3 sin 2C}{sin 2A + sin 2B + sin 2C}\rgroup"